jean is 1/3rd her brother's age plus 4 how old is jean's brother if jean is 12?
Answer by stanbon(75887)

jean is 1/3rd her brother's age plus 4 how old is jean's brother if jean is 12?
---------------
Equations:
J = (1/3)B + 4
J = 12
---
Substitute and solve for "B":
12 = (1/3)B + 4
(1/3)B = 8
B = 24 yrs (Brother's age now)
================================
Cheers,
Stan H.
========